# Beyond Sequence: Impact of Geometric Context for RNA Property Prediction

### v1

@@ C.4 Analysis of Transformer1D and Transformer1D2D
Appendix D Analysis of Noise in 3D Structures
@@
D.1 Impact of Sequence Length on 3D Structure Prediction Noise
To investigate the hypothesis that longer sequences result in greater noise in 3D structure predictions, we randomly selected a COVID and Tc-Riboswitches dataset sequence and generated structures using four state-of-the-art 3D structure prediction tools: RhoFold [45], RNAComposer [64], trRosetta [5], and SimRNA [6]. High variability among these predicted structures would indicate significant uncertainty in absolute atom positions. We quantified this noise by aligning the structures using the Kabsch algorithm [26] and computing the pairwise RMSD values, resulting in a 4x4 matrix showing structural deviations between each pair of tools (see Table 4). The observed pairwise RMSD values ranged from 16 to 45 Å for the COVID dataset and from 11 to 15 Å for the Tc-Riboswitches dataset, reflecting substantial variability and suggesting considerable noise in the 3D predictions. This level of structural inaccuracy likely contributes to the poorer performance of 3D models. However, we found that 3D models outperform 1D models for shorter sequences, such as those in the Tc-Riboswitches dataset (67 to 73 nucleotides long). This improved performance is due to the lower noise in 3D predictions for shorter sequences, a phenomenon supported by previous studies [32, 37] and also exhibited by the comparatively smaller RMSD values reported in Table 4 for Tc-Riboswitches dataset. The reduced complexity of shorter sequences allows 3D models to capture structural details more accurately, thereby enhancing performance and validating that accurate 3D structure encoding can outperform 1D models.
@@ D.1 Impact of Sequence Length on 3D Structure Prediction Noise
D.2 Impact of different 3D prediction tools
In this section, we demonstrate the significant differences in 3D structures predicted by various tools. We compare the 3D structure obtained from RhoFold [45], which serves as our default method, with those predicted by RNAComposer [64], trRosetta [5], and SimRNA [6]. Each structure is visualized side by side with RhoFold in Figure 11 to facilitate a more intuitive comparison. As observed, these structures predicted by each tool vary considerably.
@@ Appendix E Reproduction
E.1 Training details
All experiments were conducted on a single NVIDIA A100 GPU. For each baseline, hyperparameters were optimized using Optuna [1], restricting the search to models with fewer than 10 million parameters that fit within the memory constraints of an 80GB NVIDIA A100 GPU. Most baseline implementations were sourced from PyTorch Geometric [13]. The Transformer1D model was adapted to Transformer1D2D as detailed in the paper. For EGNN, we utilized the authors’ implementation [40], and for SchNet, the implementation from [25] was used.
E.2 Hyperparameters
@@
is a standard Transformer architecture for RNA sequence processing. It includes an embedding layer to convert input tokens into dense vectors, positional encoding (PE) to retain sequence order, and a multi-layer Transformer encoder to capture complex dependencies within the sequence. There are some hyperparameters from the original transformer [54]. nhead, which defines the number of attention heads in each Transformer layer; num_encoder_layers, which controls the number of encoder layers in the Transformer; d_model, which determines the dimensionality of the embeddings and the model; dim_feedforward, which sets the dimensionality of the feedforward network model. To shrink the search space, we set d_model and dim_feedforward as the same with a new hyperparameter hidden.
@@
For SchNet, the unique hyperparameters include num_filters, which refers to the number of filters used in convolutional layers, and num_gaussians, which indicates the number of Gaussian functions used for radial filters. For a more detailed explanation of these hyperparameters, please refer to [42].
@@
Appendix F Additional Results
@@
F.1 Impact of data availability
@@
F.2 Impact of partial labeling
@@
F.3 Robustness to sequencing noise

### v2

@@ F.4 Analysis of Transformer1D and Transformer1D2D
Appendix G Analysis of Noise in 3D Structures
@@
G.1 Impact of Sequence Length on 3D Structure Prediction Noise
To investigate the hypothesis that longer sequences result in greater noise in 3D structure predictions, we randomly selected a COVID and Tc-Riboswitches dataset sequence and generated structures using four state-of-the-art 3D structure prediction tools: RhoFold (Shen et al., 2022), RNAComposer (Xu et al., 2014), trRosetta (Baek et al., 2024), and SimRNA (Boniecki et al., 2016). High variability among these predicted structures would indicate significant uncertainty in absolute atom positions. We quantified this noise by aligning the structures using the Kabsch algorithm (Kabsch, 1976) and computing the pairwise RMSD values, resulting in a 4x4 matrix showing structural deviations between each pair of tools (see Table 4). The observed pairwise RMSD values ranged from 16 to 45 Å for the COVID dataset and from 11 to 15 Å for the Tc-Riboswitches dataset, reflecting substantial variability and suggesting considerable noise in the 3D predictions. This level of structural inaccuracy likely contributes to the poorer performance of 3D models. However, we found that 3D models outperform 1D models for shorter sequences, such as those in the Tc-Riboswitches dataset (67 to 73 nucleotides long). This improved performance is due to the lower noise in 3D predictions for shorter sequences, a phenomenon supported by previous studies (Nithin et al., 2024; Ponce-Salvatierra et al., 2019) and also exhibited by the comparatively smaller RMSD values reported in Table 4 for Tc-Riboswitches dataset. The reduced complexity of shorter sequences allows 3D models to capture structural details more accurately, thereby enhancing performance and validating that accurate 3D structure encoding can outperform 1D models.
@@ G.1 Impact of Sequence Length on 3D Structure Prediction Noise
G.2 Impact of different 3D prediction tools
In this section, we demonstrate the significant differences in 3D structures predicted by various tools. We compare the 3D structure obtained from RhoFold (Shen et al., 2022), which serves as our default method, with those predicted by RNAComposer (Xu et al., 2014), trRosetta (Baek et al., 2024), and SimRNA (Boniecki et al., 2016). Each structure is visualized side by side with RhoFold in Figure 11 to facilitate a more intuitive comparison. As observed, these structures predicted by each tool vary considerably.
@@ Appendix H Reproduction
H.1 Training details
All experiments were conducted on a single NVIDIA A100 GPU. For each baseline, hyperparameters were optimized using Optuna (Akiba et al., 2019), restricting the search to models with fewer than 10 million parameters that fit within the memory constraints of an 80GB NVIDIA A100 GPU. Most baseline implementations were sourced from PyTorch Geometric (Fey & Lenssen, 2019). The Transformer1D model was adapted to Transformer1D2D as detailed in the paper. For EGNN, we utilized the authors’ implementation (Satorras et al., 2021), and for SchNet, the implementation from (Joshi et al., 2023) was used.
H.2 Hyperparameters
@@
is a standard Transformer architecture for RNA sequence processing. It includes an embedding layer to convert input tokens into dense vectors, positional encoding (PE) to retain sequence order and a multi-layer Transformer encoder to capture complex dependencies within the sequence. There are some hyperparameters from the original transformer (Vaswani et al., 2017). nhead, which defines the number of attention heads in each Transformer layer; num_encoder_layers, which controls the number of encoder layers in the Transformer; d_model, which determines the dimensionality of the embeddings and the model; dim_feedforward, which sets the dimensionality of the feedforward network model. To shrink the search space, we set d_model and dim_feedforward as the same with a new hyperparameter hidden.
@@
For SchNet, the unique hyperparameters include num_filters, which refers to the number of filters used in convolutional layers, and num_gaussians, which indicates the number of Gaussian functions used for radial filters. For a more detailed explanation of these hyperparameters, please refer to (Schütt et al., 2017).
@@
Appendix I Additional Results
@@
I.1 Impact of data availability
@@
I.2 Impact of partial labeling
@@
FastEGNN 1.223 $\pm$ 0.008 0.929 $\pm$ 0.008 0.860 $\pm$ 0.004 0.837 $\pm$ 0.004 0.753 $\pm$0.015
I.3 Robustness to sequencing noise